OpenStudy (anonymous):

triple integral (cylindrical) 0 to 2pi, 0 to 3, 0 to Z/3 of r^3 dr dz dtheta; I have did this several times but I cannot get the answer of 3pi/10 some one catch my mistake?

OpenStudy (zarkon):

"some one catch my mistake?" show what you have done and I'll point out any errors

OpenStudy (anonymous):

int 0 to z/3 of r^3 dr= 1/4r^4 evaluated from 0 to z/3 = z^4/324 correct?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

int 0 to 3 of z^4/324=1/5z^5 evaluated from 0 to 3 (1/324 the constant)=243/324 or 3/4 correct?

OpenStudy (zarkon):

I get 3/20

OpenStudy (anonymous):

1/5 z^5 (1/324) yes?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

crap....i forgot to multiply by 1/5

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

nice catch thanks

OpenStudy (zarkon):

np